# Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

### v1

@@ 1 Introduction
Convolutional Neural Networks (CNN) have revolutionized the performance of image analysis tasks including image classification, semantic segmentation, object detection, and image and video synthesis [33, 7, 9, 31]. At the time of writing, CNNs continue to play a domenent role as image feature encoders for state-of-the-art techniques including several prominent Vision-Language Models (VLMs) [24, 20] as well as diffusion models for image generation [27, 32]. Nevertheless, the extraordinary complexity of CNNs has coined the nick-name of black-box, that learns an uninterpretable and high-dimensional feature representation. This work contributes toward improving our understanding of the learned representation by measuring the statistical characteristics of deep CNN features through high-dimensional statistical analysis techniques including copula analysis, and the Method of Orthogonal Moments (MOM) to obtain the Generalized Characteristic Function (GCF). Our approach is unique because it provides a novel and interpretable probability density estimate of CNN features, without making any rigid parametric assumptions that may be unjustified, and without altering the CNN feature representation which may reduce accuracy. Furthermore, as an empirical technique, we can gain greater insight by plotting and analyzing the marginal and interdependence components of the feature copula density PDF which can lead to improved understanding of the statistical behavior of the native CNN feature space. We want to measure and observe the probability density of deep features for popular CNN architectures, as we believe that this will lay the groundwork for future feature-density analysis methods that can identify stronger parametric assumptions, thereby leading to improved understanding of the feature space as well as improved probability density estimation of the distribution of deep CNN features.
@@
Measuring the probability density of deep CNN features has many practical applications including Out-of-Distribution detection [19, 21, 35, 13, 18], adversarial detection [19, 18], domain generalization [3], and federated learning [30]. It is anticipated that if one knows the density distribution of deep features, that it would be possible to use outlier detection to statistically distinguish legitimate inlier data from anomalous outlier data, thereby providing a greater level of model robustness to data that is out of model scope [8, 19, 21]. Furthermore, methods that can model both the conditional as well as the joint distribution meet the technical definition of generative methods rather than purely discrimination methods [14, 2, 15, 25]. Generative methods have many theoretical advantages because they enable complex inferencing tasks regarding the distribution of unlabeled samples in addition to performing basic discrimination tasks.
@@
Several works have attempted to construct deep generative architectures with explicit feature representations that enable exact or approximate PDF estimates for the purposes of variational inference [14, 2, 15, 25]. Such techniques include variational auto-encoders [14], as well as autoregressive and normalizing flows [2, 15, 25]. Nevertheless, open questions remain regarding the applicability of deep generative PDFs for anomaly detection for several reasons including the often unaccounted for effects of reparameterization [18] and class-imbalance [13], as well as questions regarding the empirical validity of the underlying ’typical set’ hypothesis [34] stating that there is a discernible difference between the statistical properties of typical valued in-distribution and out-of-distribution features [19, 21, 35]. We believe that further empirical analysis, including alternate ways of obtaining density estimates will be valuable toward studying these observed phenomena.
We take a purely empirical approach toward accurate estimation of deep CNN feature density. An important contribution of our approach is that we do so without unverified parametric or linearity assumptions, and without altering the CNN architecture or feature representation which may reduce accuracy. Several prior works have assumed the CNN features follow a Multi-Variate Gaussian (MVG) [22, 35, 19, 26, 35]. But the motivation for MVG traces back to a handful of histogram plots of the penultimate critic features in the supplemental materials of Lee et al. [19].
@@
There has never been a verification that MVG is suitable to the deeper intermediate features (non-penultimate), which is especially important because in most architectures, deep features undergo ReLU activation which deactivates (zeros out) any negative valued features, thereby enforcing non-negativity. Max-pooling moreover retains the strongest positive features, throwing away weaker less-positive detections. Thus it is important to look at and measure the distribution of non-negative features as these are the only features that have an impact on subsequent activations. We observe that these non-negative deep features are highly non-Gaussian, and increasingly approximate an exponential distribution with increasing network depth. We furthermore show that at adequate network depth, many deep features exhibit non-linear dependence, with typical feature values showing statistical independence for typical (non-extreme) values, but strong statistical dependence of extreme values. In other words, we observe that two deep features may be uncorrelated within their typical ranges, but if one feature is extremely large or small, then other features may show strong correlation or anti-correlation in their extreme values.
@@
This work also contributes to an an active and growing body of ongoing empirical work toward improved CNN feature understanding [1, 3, 6, 11, 23, 28]. Several recent works have attempted to visualize CNN features, in order to determine under what circumstances a model may learn or retain features that are spurious (correlated) versus invariant (causal) to semantic meaning [3, 23, 28]. We do not distinguish spurious versus invariant features in this work, however we do observe strong correspondence between extreme-valued features, suggesting that many deep features within intermediate representations may be learning correlated (or anti-correlated) semantic concepts with a statistical backdrop of uncorrelated detection noise. Allen-Zhu and Li [1] propose a multi-view hypothesis that intermediate deep CNN features may learn to detect specific views of target objects. Based on our results, we propose a follow-on hypothesis which we coin the exponential view hypothesis. That is, at adequate network depth, the distribution of feature detections for any given semantic target view follows exponential distribution. That is, prevalent small-value features correspond to a lack of detection of their respective target view, whereas rare large-value features correspond to a strong detection of the target view. As the actual target often corresponds to only a small portion of the image frame, the strong-valued detections are statistically rare occurrences but represent important computer vision signal, whereas the lack of detection, although more statistically prevalent, is primarily a noise signal. Our exponential view hypothesis offers a new perspective on the possible empirical shortcomings of the typical set hypothesis. It logically follows that the typical set may be accidentally eliminating the most important computer vision signal by excluding the extreme-value features from consideration in the typical set. We discuss this hypothesis further in the discussion section.
@@ 2.1 Copula Analysis
$\left(Y_{1},\;Y_{2},\;\dots\;,\;Y_{D}\right)\quad=\quad\left(F_{1}(X_{1}),\;F_{2}(X_{2}),\;\dots\;,F_{N}(X_{D})\right)$ (1)
@@ 2.3 Generalized Characteristic Function
The GCF refers to the straightforward extension of this technique to non-Fourier orthogonal moments. Analogously, one can recover the full copula density using an discrete inverse Fourier-like transform for any orthogonal basis set. Nevertheless, it is highly desirable to select basis functions that exhibit the following properties,
@@ 2.4 Normalized Legendre Polynomials
The Legendre Polynomials Fig 1(b,c) are a set of real-valued orthogonal basis functions over the target interval $(-1,1)$ with several desirable properties. Unlike the Chebyshev polynomials Fig 1(d), the Legendre polynomials Fig 1(c) exhibit zero integral over the interval $(-1,1)$, except trivially for the constant polynomial $P_{0}$. This zero-integral property is highly-desirable for our resultant GCF. The polynomials can be generated efficiently using Bonnet’s recurrence as follows,
@@
The Legendre polynomials in this form do not exhibit unit-length $L_{2}$ norm over the interval $(-1,1)$, as such, we propose normalizing the Legendre polynomials based on their $L_{2}$ norm in order to obtain unit-length orthogonal moments as in Fig 1(b). This normalized form is obtained as follows,
@@ 2.4.1 Real-valued Fourier Series
The real-valued Fourier basis functions in this form are shown in Fig 1(a). One can see that these basis functions correspond analogously one-to-one with the Legendre and Chebyshev polynomials, with $\phi_{t}$ exhibiting exactly $t$ roots over the target interval $(-1,1)$. This form of the Fourier series also exhibits our desired properties, as the non-constant basis functions have zero integral over the target interval $(-1,1)$. Moreover, for all basis functions the $L_{2}$ norm over the target interval is exactly equal to $1$ when presented in this form.
@@ 4.1 Analysis of Copula Marginals
As we see see from figure 5, the Gamma and Weibull distributions show a significantly better fit to the feature marginals than the Uniform and Gaussian distributions. The Exponential distribution can be seen as having performance in-between the Gaussian and Weibull distribution. Notably, we observe that the exponential distribution increasingly approximates the observed features as we look at deeper network depth. It is important to note that the Gamma and Weibull distributions generalize the exponential distribution, and thus it is not possible for the Exponential to achieve a statistically better fit than the Gamma and Weibull distributions. Nevertheless, the exponential distribution is highly interpretable and exhibits a very good fit to the non-zero features that significantly outperforms the Gaussian and uniform distributions, while coming close to the Gamma and Weibull goodness of fit particularly for the deeper layers.
@@ 4.2 Analysis of Copula Interdependence
Figure 6 shows the interdependence term of the copula interdependence density for resnet18 over Imagenette2 for select pairwise features. Additional plots for other feature pairs, architectures and datasets are available in supplemental materials. Figure 6 also shows the reconstructed interdependence density function using three different methods, the GCF with normalized Legendre polynomials (left), the GCF with real-valued Fourier series (middle) and 2D histograms as a control (right). We see that all three functions appear to show very good agreement with very similar pairwise density plots for all of the feature pairs shown in Figure 6.
We observe that the early layers show much more complicated interdependence with greater variation, whereas later layers tend to show a similar structure for most feature pairs that we observed. Similar to our analysis of the marginals, we believe that the variation in the pairwise interdependence seen in the shallow layers (layer 0 and layer 1) is very likely due to the contamination of the input pixel and texture distribution into the shallow features. As these are early layers in the network, the network architecture is unable to fully remove the contribution of the initial pixel distribution. The later layers however show very similar pairwise interdependence. We observe a very interesting phenomenon, in which the deep features in layers 2,3 and 4 (rows 3-5 in fig 6 appear to be uncorrelated for typical values, but show a strong statistical dependence for extreme valued features. This strong statistical dependence can be observed in figure 6 (bottom three rows) by the prevalence of very high density in the upper right corner of the interdependence plot (yellow), with much lower density along the top and right edges of the plot (blue). The apparent statistical independence of the terms over typical values (green) is also clearly apparent in the pairwise interdependence plots. We also observed (shown in supplemental materials) the prevalence of deep features showing a strong anti-correlation of the extreme values, with high density (yellow) along the top and right, but low density in the upper right hand corner (blue), again with independence for the typical valued features (green). This prevalence of uncorrelated features, except in the event of extreme-valued detections, is a new observation with profound implications regarding the distribution of deep CNN features, we discuss this result further in our discussion section.
@@
In order to evaluate the goodness-of-fit of the copula independence term, we perform a supervised evaluation using cross entropy loss. Our task is to fit the copula interdependence to the training features using our method, and then we evaluate how well this describes the interdependence of the test features. For this experiment, we calculated the goodness of fit for resnet18 features for random groups of four features at a time for the imagenette2 dataset. We performed 30 rounds per experiment, and calculated the copula interdependence using 3 methods, Legendre GCF, Fourier GCF, and Histograms. The interdependence of four randomly selected features within the same layer is calculated, and compared with the sample interdependence of the test features using cross entropy loss as the evaluation criteria. We observe that the GCF methods significantly outperform the Histogram methods for Layers 2,3 and 4, with the only exception being the Legendre GCF in Layer 3. The Histogram approach is particularly sensitive to the CoD, because with $D$ dimensions and $B$ bins per dimension, the average number of samples per bin is $N/B^{D}$. Due to max-pooling, the deeper layers exhibit fewer samples relative to the shallow layers, and one would expect the GCF methods to outperform the histogram methods particularly in the situations of higher dimensionality as well as smaller sample sizes, which is consistent with our result in table 2.

### v2

@@ 1 Introduction
Convolutional Neural Networks (CNN) have revolutionized the performance of image analysis tasks including image classification, semantic segmentation, object detection, and image and video synthesis [34, 7, 9, 32]. At the time of writing, CNNs continue to play a domenent role as image feature encoders for state-of-the-art techniques including several prominent Vision-Language Models (VLMs) [22, 25, 20] as well as diffusion models for image generation [28, 33]. Nevertheless, the extraordinary complexity of CNNs has coined the nick-name of black-box, that learns an uninterpretable and high-dimensional feature representation. This work contributes toward improving our understanding of the learned representation by measuring the statistical characteristics of deep CNN features through high-dimensional statistical analysis techniques including copula analysis, and the Method of Orthogonal Moments (MOM) to obtain the Generalized Characteristic Function (GCF). Our approach is unique because it provides a novel and interpretable probability density estimate of CNN features, without making any rigid parametric assumptions that may be unjustified, and without altering the CNN feature representation which may reduce accuracy. Furthermore, as an empirical technique, we can gain greater insight by plotting and analyzing the marginal and interdependence components of the feature copula density PDF which can lead to improved understanding of the statistical behavior of the native CNN feature space. We want to measure and observe the probability density of deep features for popular CNN architectures, as we believe that this will lay the groundwork for future feature-density analysis methods that can identify stronger parametric assumptions, thereby leading to improved understanding of the feature space as well as improved probability density estimation of the distribution of deep CNN features.
@@
Measuring the probability density of deep CNN features has many practical applications including Out-of-Distribution detection [19, 21, 36, 13, 18], adversarial detection [19, 18], domain generalization [3], federated learning [31], and semi-supervised learning [23]. It is anticipated that if one knows the density distribution of deep features, that it would be possible to use outlier detection to statistically distinguish legitimate inlier data from anomalous outlier data, thereby providing a greater level of model robustness to data that is out of model scope [8, 19, 21]. Furthermore, methods that can model both the conditional as well as the joint distribution meet the technical definition of generative methods rather than purely discrimination methods [14, 2, 15, 26]. Generative methods have many theoretical advantages because they enable complex inferencing tasks regarding the distribution of unlabeled samples in addition to performing basic discrimination tasks.
@@
Several works have attempted to construct deep generative architectures with explicit feature representations that enable exact or approximate PDF estimates [14, 2, 15, 26]. Such techniques include variational auto-encoders [14], as well as autoregressive and normalizing flows [2, 15, 26]. Nevertheless, open questions remain regarding the applicability of deep generative PDFs for anomaly detection for several reasons including the often unaccounted for effects of reparameterization [18] and class-imbalance [13], as well as questions regarding the empirical validity of the underlying ’typical set’ hypothesis [35] stating that there is a discernible difference between the statistical properties of typical valued in-distribution and out-of-distribution features [19, 21, 36]. We believe that further empirical analysis, including alternate ways of obtaining density estimates will be valuable toward studying these observed phenomena.
@@
We take a purely empirical approach toward accurate estimation of deep CNN feature density. An important contribution of our approach is that we do so without unverified parametric or linearity assumptions, and without altering the CNN architecture or feature representation which may reduce accuracy. Several prior works have assumed the CNN features follow a Multi-Variate Gaussian (MVG) [23, 36, 19, 27, 36]. But the motivation for MVG traces back to a handful of histogram plots of the penultimate critic features in the supplemental materials of Lee et al. [19]. There has never been a verification that MVG is suitable to the deeper intermediate features (non-penultimate), which is especially important because in most architectures, deep features undergo ReLU activation which deactivates (zeros out) any negative valued features, thereby enforcing non-negativity. Max-pooling moreover retains the strongest positive features, throwing away weaker less-positive detections. Thus it is important to look at and measure the distribution of non-negative features as these are the only features that have an impact on subsequent activations. We observe that these non-negative deep features are highly non-Gaussian, and increasingly approximate an exponential distribution with increasing network depth. We furthermore observe that at adequate network depth, many deep features exhibit non-linear dependence, with typical feature values showing statistical independence for typical (non-extreme) values, but strong statistical dependence of extreme values. In other words, we observe that two deep features may be uncorrelated within their typical ranges, but if one feature is extremely large or small, then other features may show strong correlation or anti-correlation in their extreme values.
@@
This work also contributes to an active and growing body of ongoing empirical work toward improved CNN feature understanding [1, 3, 6, 11, 24, 29]. Several recent works have attempted to visualize CNN features, in order to determine under what circumstances a model may learn or retain features that are spurious (correlated) versus invariant (causal) to semantic meaning [3, 24, 29]. We do not distinguish spurious versus invariant features in this work, however we do observe strong correspondence between extreme-valued features, suggesting that many deep features within intermediate representations may be learning correlated (or anti-correlated) semantic concepts with a statistical backdrop of uncorrelated detection noise. Allen-Zhu and Li [1] propose a multi-view hypothesis that intermediate deep CNN features may learn to detect specific views of target objects. Based on our results, we propose a follow-on hypothesis which we coin the exponential view hypothesis. That is, at adequate network depth, the distribution of feature detections for any given semantic target view follows exponential distribution. That is, prevalent small-value features correspond to a lack of detection of their respective target view, whereas rare large-value features correspond to a strong detection of the target view. As the actual target often corresponds to only a small portion of the image frame, the strong-valued detections are statistically rare occurrences but represent important computer vision signal, whereas the lack of detection, although more statistically prevalent, is primarily a noise signal. Our exponential view hypothesis offers a new perspective on the possible empirical shortcomings of the typical set hypothesis. It logically follows that the typical set may be accidentally eliminating the most important computer vision signal by excluding the extreme-value features from consideration in the typical set. We discuss this hypothesis further in the discussion section.
@@ 2.1 Copula Analysis
$\left(Y_{1},\;Y_{2},\;\dots\;,\;Y_{D}\right)\quad=\quad\left(F_{1}(X_{1}),\;F_{2}(X_{2}),\;\dots\;,F_{D}(X_{D})\right)$ (1)
@@ 2.3 Generalized Characteristic Function
The GCF refers to the straightforward extension of this technique to non-Fourier orthogonal moments. Analogously, one can recover the full copula density using a discrete inverse Fourier-like transform for any orthogonal basis set. Nevertheless, it is highly desirable to select basis functions that exhibit the following properties,
@@ 2.4 Normalized Legendre Polynomials
The Legendre Polynomials figure 1(b,c) are a set of real-valued orthogonal basis functions over the target interval $(-1,1)$ with several desirable properties. Unlike the Chebyshev polynomials figure 1(d), the Legendre polynomials figure 1(c) exhibit zero integral over the interval $(-1,1)$, except trivially for the constant polynomial $P_{0}$. This zero-integral property is highly-desirable for our resultant GCF. The polynomials can be generated efficiently using Bonnet’s recurrence as follows,
@@
The Legendre polynomials in this form do not exhibit unit-length $L_{2}$ norm over the interval $(-1,1)$, as such, we propose normalizing the Legendre polynomials based on their $L_{2}$ norm in order to obtain unit-length orthogonal moments as in figure 1(b). This normalized form is obtained as follows,
@@ 2.4.1 Real-valued Fourier Series
The real-valued Fourier basis functions in this form are shown in figure 1(a). One can see that these basis functions correspond analogously one-to-one with the Legendre and Chebyshev polynomials, with $\phi_{t}$ exhibiting exactly $t$ roots over the target interval $(-1,1)$. This form of the Fourier series also exhibits our desired properties, as the non-constant basis functions have zero integral over the target interval $(-1,1)$. Moreover, for all basis functions the $L_{2}$ norm over the target interval is exactly equal to $1$ when presented in this form.
@@ 4.1 Analysis of Copula Marginals
As we see from figure 5, the gamma and Weibull distributions show a significantly better fit to the feature marginals than the uniform and Gaussian distributions. The exponential distribution can be seen as having performance in-between the Gaussian and Weibull distribution. Notably, we observe that the exponential distribution increasingly approximates the observed features as we look at deeper network depth. It is important to note that the gamma and Weibull distributions generalize the exponential distribution, and thus it is not possible for the exponential to achieve a statistically better fit than the gamma and Weibull distributions. Nevertheless, the exponential distribution is highly interpretable and exhibits a very good fit to the non-zero features that significantly outperforms the Gaussian and uniform distributions, while coming close to the gamma and Weibull goodness of fit particularly for the deeper layers.
@@ 4.2 Analysis of Copula Interdependence
Figure 6 shows the interdependence term of the copula interdependence density for resnet18 over imagenette2 for select pairwise features. Additional plots for other architectures and datasets are available in supplemental materials. Figure 6 also shows the reconstructed interdependence density function using three different methods, the GCF with normalized Legendre polynomials (left), the GCF with real-valued Fourier series (middle) and 2D histograms as a control (right). We see that all three functions appear to show very good agreement with very similar pairwise density plots for all of the feature pairs shown in figure 6.
We observe that the early layers show much more complicated interdependence with greater variation, whereas later layers tend to show a similar structure for most feature pairs that we observed. Similar to our analysis of the marginals, we believe that the variation in the pairwise interdependence seen in the shallow layers (layer 0 and layer 1) is very likely due to the contamination of the input pixel and texture distribution into the shallow features. As these are early layers in the network, the network architecture is unable to fully remove the contribution of the initial pixel distribution. The later layers however show very similar pairwise interdependence. We observe a very interesting phenomenon, in which the deep features in layers 2,3 and 4 (rows 3-5 in figure 6 appear to be uncorrelated for typical values, but show a strong statistical dependence for extreme valued features. This strong statistical dependence can be observed in figure 6 (bottom three rows) by the prevalence of very high density in the upper right corner of the interdependence plot (yellow), with much lower density along the top and right edges of the plot (blue). The apparent statistical independence of the terms over typical values (green) is also clearly apparent in the pairwise interdependence plots. We also observed (shown in supplemental materials) the prevalence of deep features showing a strong anti-correlation of the extreme values, with high density (yellow) along the top left and bottom right, but low density in the upper right hand corner (blue), again with independence for the typical valued features (green). This prevalence of uncorrelated features, except in the event of extreme-valued detections, is a new observation with profound implications regarding the distribution of deep CNN features, we discuss this result further in our discussion section.
@@
Table 3: Goodness of fit of copula interdependence term for random subsets of four features on cifar10 with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.
Legendre Fourier Histogram
Layer 0 2.2306 (2.2174, 2.2438) 2.2359 (2.2229, 2.2490) 2.1850 (2.1712, 2.1989)
Layer 1 2.5844 (2.5830, 2.5859) 2.5812 (2.5797, 2.5827) 2.5578 (2.5548, 2.5608)
Layer 2 2.6834 (2.6828, 2.6841) 2.6852 (2.6846, 2.6858) 2.6865 (2.6858, 2.6873)
Layer 3 2.7442 (2.7440, 2.7444) 2.7461 (2.7459, 2.7463) 2.7493 (2.7491, 2.7495)
Layer 4 2.1568 (2.1502, 2.1634) 2.1268 (2.1204, 2.1332) 2.0399 (2.0339, 2.0459)
Table 4: Goodness of fit of copula interdependence term for random subsets of four features on cifar100 with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.
Legendre Fourier Histogram
Layer 0 2.2020 (2.1812, 2.2227) 2.2085 (2.1883, 2.2288) 2.0857 (2.0644, 2.1070)
Layer 1 2.5588 (2.5571, 2.5605) 2.5600 (2.5583, 2.5618) 2.5242 (2.5219, 2.5264)
Layer 2 2.6794 (2.6788, 2.6800) 2.6806 (2.6800, 2.6812) 2.6829 (2.6822, 2.6835)
Layer 3 2.7414 (2.7414, 2.7414) 2.7415 (2.7415, 2.7415) 2.7449 (2.7449, 2.7450)
Layer 4 2.7200 (2.7198, 2.7202) 2.7220 (2.7218, 2.7222) 2.7303 (2.7301, 2.7304)
Table 5: Goodness of fit of copula interdependence term for random subsets of four features on mnist with resnet18. Evaluation criteria is cross entropy loss, $1\sigma$ confidence intervals are shown in parentheses. Bold shows significantly best method for each layer task.
Legendre Fourier Histogram
Layer 0 -3.3750 (-3.6474, -3.1025) -2.7168 (-2.8623, -2.5714) -3.6498 (-3.9232, -3.3764)
Layer 1 1.3320 (1.2504, 1.4137) 1.2019 (1.1019, 1.3019) 0.5817 (0.3903, 0.7730)
Layer 2 2.1437 (2.1295, 2.1579) 2.1298 (2.1136, 2.1460) 1.8885 (1.8526, 1.9244)
Layer 3 2.6451 (2.6444, 2.6458) 2.6551 (2.6545, 2.6557) 2.6481 (2.6472, 2.6489)
Layer 4 2.1646 (2.1467, 2.1824) 2.1572 (2.1376, 2.1768) 2.0449 (2.0147, 2.0751)
In order to evaluate the goodness-of-fit of the copula independence term, we perform a supervised evaluation using cross entropy loss. Our task is to fit the copula interdependence to the training features using our method, and then we evaluate how well this describes the interdependence of the test features. For this experiment, we calculated the goodness of fit for resnet18 features for random groups of four features at a time for the imagenette2, cifar10, cifar100 and mnist dataset. We performed 30 rounds per experiment, and calculated the copula interdependence using 3 methods, Legendre GCF, Fourier GCF, and Histograms. The interdependence of four randomly selected features within the same layer is calculated, and compared with the sample interdependence of the test features using cross entropy loss as the evaluation criteria.
For resnet18 on cifar10, Legendre GCF works better for layers 2 and 3; however, for the latest layer, the Histogram method performs the best. Table 4 demonstrates that on cifar100, Legendre GCF performs well across all layers except layer 1, where the Histogram approach outperforms due to pixels still being contaminated by the input image. On mnist (table 5), the Histogram method is the best for all layers except layer 3, where Legendre GCF outperforms. This is because mnist is essentially a binary dataset with most pixels being either black or white. This binary input pixel distribution contaminates the deep feature distributions, causing a discreet quantization which can be better modeled by the discreet Histogram approach. For the other datasets (imagenette2, cifar10, cifar100) with more continuous input pixel distributions, we observe in tables 2-4 that the GCF methods significantly outperform the Histogram methods for layers 2,3 and 4, with the two exceptions being layer 3 for imagenette2 dataset, and layer 4 for the cifar10 dataset. The Histogram approach is particularly sensitive to the CoD, because with $D$ dimensions and $B$ bins per dimension, the average number of samples per bin is $N/B^{D}$. Due to max-pooling, the deeper layers exhibit fewer samples relative to the shallow layers, and one would expect the GCF methods to outperform the histogram methods particularly in the situations of higher dimensionality as well as smaller sample sizes, which is consistent with our results in tables 2-4.